Selected data for A, inc. and B corp, two companies in the same industry, are presented below : Sorenta, Inc New world corp Sales $50 000 80 000 Cost of goods sold 30 000 50 000 Average inventory balance 5000 5000 Based on this data which of the following statements is true ? a- A has lower storage costs and a lower investment in inventory than B b- A has a lower gross profit ratio than c- B has a higher net income than A d- B sells its inventory faster than A
Answer #1

Answer:

Correct answer is:

d- B sells its inventory faster than A

Explanation:

Inventory turnover of A = Cost of goods sold / Average inventory = 30000 / 5000 = 6

Inventory turnover of B = 50000 / 5000 = 10

Inventory turnover of B is higher than A; hence B sells its inventory faster than A

As such option d is correct.

Option a is incorrect since both A and B have equal average inventory and hence both are expected to have equal inventory storage costs

Option b is incorrect since gross profit ratio of A is [((50000-30000)/50000)=] 40% which is higher than gross profit ratio of B which is [((80000-50000)/80000)=] 37.5%.

For option c, it is difficult to say  B has a higher net income than A, since data to get Net Income are not given.